Jill has $3.50 in nickels and dimes. If she has 50 coins, how many of each type of coin does she have?
Answer by checkley71(8403)

N+D=50
10D+5N=350
N=50-D OR SUBSTITUTING 50-D FOR N IN EQUATION #2 WE GET
10D+5(50-D)=350 OR 10D+250-5D=350 OR 5D=100 OR D=100/5 OR D=20 & N+20=50 OR N=30

PROOF 10*20+5*30=350 OR 200+150=350 OR 350=350